?
Asletien of Hi sbaced by seahon wit pure sadiun Car bomde FA105 43o oot) ind the madarily ot the HNos and ls absolute uncertainty Number
Answer #1

Page No: Date A volume of 29-32 :土0.oSmL of HNO3 Soluhencou CFM lo0oo) Find Malas HNO3 Solve thiu and SmatutvHNo3 with theie valu }.eqof anduwer will b lanpet が.qvtー9:od): teS. 9? (N,ν,-Ns%) ογ HNb3 MNO3 1o 5.98 looo os 991x 29.21 --> Simil lowest v alu Mslau 91 29.31 ml valumu a HLA 29 31HNo3 loop 2N1=normality of na2co3 and v1= volume of na2co3

and N2=normality of hno3 and V2= volume of hno3

N1V1=N2V2

normality = molarity * n factor

for hno3 n factor=1

for na2co3 n factor=2*M1*V1=M2V2

we have to put the values in such a way to get minimum and maximum values of molarity.

hence the formula in terms of molarity will be 2 * M1*V1=M2*V2